# LibriQuote: A Speech Dataset of Fictional Character Utterances for Expressive Zero-Shot Speech Synthesis

### v1

@@ Acknowledgements
Appendix A Technical Appendix
Further Descriptive Statistics and Comparisons
In this section, we provide additional information regarding LibriQuote and its comparison to LibriHeavy.
Quotations and Pitch Variation
Audiobook narrators typically alternate between neutral utterances of narration, and expressive utterances in the form of character quotations. Sometimes, they may even completely incarnate a character: for instance, a male speaker might raise his pitch to quote a small female child character. This pitch variation reflects in datasets such as LibriHeavy: audio utterances are typically built by taking around 30 seconds of audio, randomly segmenting at sentence boundaries. Hence, they might contain a various number of quotations, leading to very complex pitch distributions. As a result, Text-to-speech (TTS) systems trained on such audio segments might be hindered by being trained to reproduce these complex distributions without access to meaningful information, such as the fact that a raise in pitch might be favored by the impersonation of a little girl.
To illustrate this effect, we used the small subset of LibriHeavy, and performed quotation detection on each transcription. We found that 25% of audio segments contained 1 to 12 quotations. Then, we calculate pitch standard deviation using Parselmouth (Boersma and Weenink 2021) for each segment and aggregate the results per number of quotations in the segment. The results are displayed in Figure 4. We see that pitch distributions are typically more spread when the number of quotations in the segments increases. This result is supported by looking at Spearman correlations between the number of quotations in the segment and pitch standard deviations (, ). In comparison, LibriQuote-test quotations exhibit a pitch standard deviation of ( bootstrapped 95% confidence interval). Besides, LibriQuote offers 5300 hours of training data distributed among more than 3 millions utterances that are deliberately separated from neutral narration style.
Accent Representations
LibriVox is a collaborative platform where anyone can engage in recording a segment of audiobook. Hence, it contains a varied panel of speakers and accents. We illustrate this in Figure 5 by looking at LibriQuote-test quotations t-SNE projections of accent representations computed with an ECAPA-TDNN model trained on Common Voice101010https://huggingface.co/Jzuluaga/accent-id-commonaccent˙ecapa (Desplanques, Thienpondt, and Demuynck 2020; Ardila et al. 2020; Zuluaga-Gomez et al. 2023). We see a wide range of clusters, that can be interpreted as a large set of accents being represented in the test data.
Emotion Representations
We display in Table 6 the proportion of predicted emotion labels using emotion2vec-plus-base on LibriQuote-test quotations, reference narration and LibriHeavy samples. We used the base variant of emotion2vec plus rather than large as we observed that the large variant was producing collapsed representations, with a very low variance in cosine similarity values. We observe a broader coverage of predicted emotions in quotations compared to their narration references and LibriHeavy audio segments.
Details - Narrative Information
Prompting Experiments
@@ Appendix A Technical Appendix
Additionally, we tested various Large Language Models (LLMs) for the task of narrative information extraction, and also tested the extraction of other elements such as nouns and adjectives. Nouns and adjectives often occur when the utterance description is supplemented by additional information (e.g. “he said with a sigh” for nouns and “he added in a loud, sharp tone” for adjectives). Note that to ensure a high-quality filtering, we tried to maximize the precision of predictions in order to avoid potential hallucinations where some narrative information are spuriously associated to utterances. We see in Table 7 that all tested models seem to perform relatively well on verbs, but some across-model variance is manifested for adverbs. We found that Qwen2.5-72b provides the best overall performance for these two types of information, followed closely by Phi-4. In the end, we decided to use Phi-4 to extract information on the full train set, as it obtained very high precision in 10 less computation time than Qwen2.5-72b. Besides, we found that LLMs dramatically fail at extracting accurately adjectives and nouns, with very low F1 scores. We found during the annotation that extracting such information is harder, even for humans, as introducing adjectives and nouns is less grounded in linguistic rules, and authors might use a more diverse set of linguistic devices to introduce these information. Thus, we decided not to rely on the extraction of such elements.
Overall, extracting verbs and adverbs with Phi-4 (we used vLLM (Kwon et al. 2023) as the inference backend) on the full train set took around 6 hours on a single NVIDIA H100 GPU card.
| LibriQuote (Q) | LibriQuote (R) | LibriHeavy | |
@@
To build the list of speech verbs, , we started by extracting a large list of 201 potential speech verbs from the web111111https://archiewahwah.wordpress.com/speech-verbs-list/ along with their descriptions. Then, based on verb descriptions, we discarded every verb that might indicate a neutral way of speaking. The resulting verb list contains 89 speech verbs and can be found in Table 8.
| Prompts | Verbs | Adverbs | Adjectives | Nouns | |||||||||
@@
| - | 0.98 | 0.8 | 0.88 | 0.93 | 0.73 | 0.82 | 0.67 | 0.19 | 0.25 | 0.38 | 0.21 | 0.16 | |
Details - Baselines
@@
is an autoregrssive TTS system that builds upon Qwen2.5-0.5M. It also uses BiCodec, a speech codec model that transforms 16kHz sampling rate speech into two types of discrete tokens: global and semantic tokens. Global tokens are derived from representations of an ECAPA-TDNN model fine-tuned for speaker-verification, hence capturing speaker-related information such as timbre. The ECAPA-TDNN representations are mapped to a set of 32 hidden features using learned laten queries, which are then discretized using Finite Scalar Quantization (Mentzer et al. 2023). Semantic tokens are derived from representations of the 11th, 14th and 16th layers of Wav2Vec 2.0 (XLSR-53) (Pasad, Shi, and Livescu 2023). These features are then quantized using single-codebook vector quantization, using factorized codes as done in DAC (Kumar et al. 2023a). Tokens are converted back to raw audio using a Convolutional Neural Network (CNN) Decoder composed of ConvNeXt blocks. The language modeling framework is used to learn to generate semantic tokens, conditioned on the text-to-synthesis tokens and global tokens (which are prepended in the autoregressive sequence).
@@
is a non-autoregressive diffusion TTS system based on Flow Matching that does not rely on phonetic alignment nor duration predictor, operating on 24kHz sampling rate speech. During training, following the text-guided speech infilling task, speech mel spectrograms are first masked and then concatenated with text tokens. The resulting sequence is inputted to a Diffusion Transformer (DiT) (Peebles and Xie 2023b), which learns to generate the masked spectrogram region from gaussian noise using Flow Matching. During inference, the reference speech mel spectogram serves as the unmasked region, and is concatenated with the text to synthesize the input sequence. Sway sampling is used along with an Ordinary Differential Equation (ODE) solver to generate mel spectrograms that are converted back using Vocos (Siuzdak 2024).
@@
is also a non-autoregressive TTS system, that follows a two-step masked generative modeling paradigm. It leverages two speech codec models to produce semantic and acoustinc discrete units from 24kHz sampling rate speech. Semantic tokens are derived from representations of the 17th layer of W2v-BERT 2.0, that are converted to discrete units using a single codebook of size 8,192 via classic Vector Quantization (VQ). Acoustic tokens are produced following RVQGAN (Kumar et al. 2023b), and 8 codebooks of size 8,192 are used via Residual VQ (RVQ). To generate raw audio, MaskGCT follows a two-step process: first, it uses a Transformer to generate semantic tokens by iteratively predicting masked tokens conditioned on the text to synthesize, prompt tokens and previously generated semantic tokens ; second acoustic tokens are generated following a similar iterative procedure, but using an architecture similar to SoundStorm (Borsos et al. 2023), and the generative process is conditioned with the prompt acoustic tokens and previously predicted semantic tokens. Finally, predicted acoustic tokens are converted back to raw audio using Vocos. Additionally, MaskGCT employs a Flow-Matching duration predictor to predict the target speech duration during inference.
Details - Subjective Experiments
@@
For the CMOS experiments, we ask the participants the following question: How expressive is this recording? Please focus on aspects of style (timbre, emotion and prosody), and ignore the aspects of content, grammar, or audio quality. For each instance, we give participants two audios containing one human sample and one generated sample, in a randomized order. We ask the participants to rate if the first audio sounds more expressive than the second on a scale. Screenshot of the CMOS platform can found in Figure 7
For the MOS experiments, we ask the participants the following question: How natural (i.e human-sounding) is this recording? Please focus on examining the audio quality and naturalness, and ignore aspects of style (timbre, emotion, prosody). For each instance, we give participants a single audio, and ask them to rate the naturalness of the sample on an 1-5 Likert scale with 0.5 intervals. Screenshot of the MOS platform can be found in Figure 8
Hardware Information
@@
Reproducibility Checklist
We would like to address the following points in the Reproducibility Checklist:
All novel datasets introduced in this paper are included in a data appendix. LibriQuote is a speech dataset that include a large amount of audio files, and thus can not be shared straightforwardly in a Data appendix.
For the full dataset release, we plan to release alignments in JSON formats on Huggingface, where a typical LibriQuote instance will comes with start and end anchors (in absolute second position) from an original LibriLight audio file. We will also share scripts to automatically download LibriLight audio recordings and to build the full dataset given the alignment information.
For the data shared in the submission, we provide: 1) a demo page with audio samples used in the listening tests (available at https://libriquote.github.io/) and 2) the alignment information of LibriQuote-test, in case reviewers are interested in looking in details at the test-data.
@@
”whispered”: {
”id”: ”2”,
”type”: ”verb”,
”confidence”: 10
”slowly”: {
”id”: ”2”,
”type”: ”adverb”,
”confidence”: 10
@@
”went”: {
”id”: ”0”,
”type”: ”verb”,
”confidence”: 9
”laughing”: {
”id”: ”0”,
”type”: ”verb”,
”confidence”: 9

### v2

@@ 8 Ethical Considerations
We also note that LibriQuote does not provide explicit gender or accent distributions of its speakers. As a results, TTS models trained with LibriQuote are likely to exhibit gender or accent bias ; where one gender or one/multiple accents are either generated more faithfully or generated with stereotypical cues. We think researchers and/or companies using LibriQuote as training data in the purpose of creating a commercial tool should systematically assess and, if possible, mitigate such bias before providing users with an end product.
@@ 9 Acknowledgements
Appendix A LibriQuote Analysis
As mentioned in Section 3.2, we performed emotion detection with the open source Emotion2Vec-plus-base. We report the detection results in Table 5 and display t-SNE projections of the resulting utterance representations in Figure 3, and compare LibriQuote-test quotations, narrations and LibriHeavy segments. As expected, we see a much higher proportion of quotations predicted as non-neutral in LibriQuote quotations compared to LibriHeavy and narration segments.
Appendix B Narrative Information
B.1 Prompting Experiments
@@
Additionally, we tested various Large Language Models (LLMs) for the task of narrative information extraction, and also tested the extraction of other elements such as nouns and adjectives. Full results can be found in Table 6. Nouns and adjectives often occur when the utterance description is supplemented by additional information (e.g. “he said with a sigh” for nouns and “he added in a loud, sharp tone” for adjectives). Note that to ensure a high-quality filtering, we tried to maximize the precision of predictions in order to avoid potential hallucinations where some narrative information are spuriously associated to utterances. We found that Qwen2.5-72b with self-confidence (SC) provides the best overall performance for these two types of information, followed closely by Phi-4 with SC. In the end, we decided to use Phi-4 with SC to extract information on the full train set, as it obtained very high precision in 10 less computation time than Qwen2.5-72b with SC. Besides, we found that LLMs dramatically fail at extracting accurately adjectives and nouns, with very low F1 scores. We found during the annotation that extracting such information is harder, even for humans, as introducing adjectives and nouns is less grounded in linguistic rules, and authors might use a more diverse set of linguistic devices to introduce these information. Thus, we decided not to rely on the extraction of such elements.
Overall, extracting verbs and adverbs with Phi-4 with SC (we used vLLM Kwon et al. (2023) as the inference backend) on the full train set took around 6 hours on a single NVIDIA H100 GPU card.
| Quotations | Narrations | LibriHeavy | |
@@ B.2 Filtering
To build the list of speech verbs, , we started by extracting a large list of 201 potential speech verbs from the web888https://archiewahwah.wordpress.com/speech-verbs-list/ along with their descriptions. Then, based on verb descriptions, we discarded every verb that might indicate a neutral way of speaking. The resulting verb list contains 89 speech verbs and can be found in Table 7.
| Utterance | Verbs | Adverbs | Adjectives | Nouns | |||||||||
@@
| Qwen2.5-72b (SC) | - | 0.98 | 0.8 | 0.88 | 0.93 | 0.73 | 0.82 | 0.67 | 0.19 | 0.25 | 0.38 | 0.21 | 0.16 |
Appendix C SparkTTS Experiments
Our experiments described in Section 4 involved fine-tuning and training from scratch numerous variants of SparkTTS with LibriQuote. We provide details below for complete reproducibility of our experiments. Whether for fine-tuning or training from scratch, we transform each quotation utterance in the sequence where are the text tokens, are SparkTTS global tokens (with a fixed number of 32 tokens) and are semantic tokens, with varying length depending on the utterance length. Each training experiments involve optimizing the following language modeling task:
in other words, we optimize the Language Model backbone solely on semantic tokens.
C.1 Fine-Tuning
For both fine-tuning experiments, we employ Adam optimizer Kingma and Ba (2017) with peak learning rate of 1e-5. We train each model for 3 epochs on their respective data split ( and . A cosine schedule is used with 10000 warmup steps for the model fine-tuned on () and 5000 steps when using (). Note that contains less utterances than . In each scenario, we form batches of 32 utterances. During inference, we compute global tokens using reference narration utterances, and prepend the text tokens and global tokens before starting the generation, following SparkTTS setup, yielding the following input sequence where are global tokens computed from the reference narration.
C.2 Training from Scratch
Here, our experiments include two variants. The first variant is similar to SparkTTS and uses the text to synthesize as the text-condition, with the unchanged sequence and training objective . The second variant leverages contextual information around quotations by replacing the text condition with the sequence where and are the text tokens from the left and right context respectively, are the tokens from the text to synthesize, and and are special delimiter tokens. Note that we also replace each quotation appearing in and with a special token. Therefore, most of the contextual information can be seen as narrative details. We use a context of one paragraph for both and . We train each variant for 10 epochs with Adam optimizer, a peak learning rate of 3e-5 and a batch size of 32. A cosine schedule is used with 20K warmup steps.
Appendix D F5-TTS Experiments
Our experiments also involve fine-tuning F5-TTS on LibriQuote’s expressive subset . We use the official fine-tuning script from the f5-tts repository999https://github.com/SWivid/F5-TTS/blob/main/src/f5_tts/train/train.py, starting with the official V1-Base checkpoint. We use a batch size of 38400 frames per GPU (with 2 H100 GPUs) and a maximum of 32 sequences per batch. The learning rate is set to 7.5e-5, and we use 20000 warmup steps, training for a total of 3 epochs.
Appendix E Contextual Metrics
With recent breakthroughs in Large Language Models (LLMs) and Large Audio Language Models (LALMs) understanding capabilities, automated evaluations of medium-sized dedicated test samples with LLM-as-a-judge Gu et al. (2025) or LALM-as-a-judge Manku et al. (2025) have become increasingly popular. These approaches delegate the usual human inspection of some test samples to LLMs or LALMs to provide an aggregated response, which was often proved to correlate strongly with human judges Gu et al. (2025); Manku et al. (2025). While human judgements would be ideal to evaluate to which extent a quotation sample convey the appropriate emotion and/or voice intent given a particular book context, subjective experiments at scale are expensive in terms of time and costs, and require dedicated knowledge to construct and analyse. LALM-as-a-judge is thus used as a replacement, noting that a model individual conclusions are often themselves biased, and that only aggregated scores are usually worth exploring.
We follow Manku et al. (2025) and use Gemini-2.5 Pro as the model judge. We hypothesize that the combined ability of LALMs to perform audio and text reasoning is valuable in our particular setup, where text reasoning is suited for understanding the right intent to convey for a particular quotation, and audio reasoning for understanding if a sample appropriately exhibits this intent. We propose two metrics that are built by modifying prompts provided in Manku et al. (2025): ContextMOS and Win-Rate. ContextMOS takes a single audio sample, a quotation text and its context, and prompts Gemini-2.5 Pro to assign a score between 1 and 5, where 1 indicates no alignment between the sample intent and the intents stemming from the context, while 5 indicates complete alignment and naturalness in conveying this intent. In contrast, Win-Rate takes a ground-truth audio sample and a synthesized sample, a quotation text and its context, and chose which audio sample conveys the right intent the most faithfully. Note that we allow ties, and define the Win-Rate to of a system relative to the ground-truth as the following:
Appendix F Baselines
@@
is an autoregrssive TTS system that builds upon Qwen2.5-0.5M. It also uses BiCodec, a speech codec model that transforms 16kHz sampling rate speech into two types of discrete tokens: global and semantic tokens. Global tokens are derived from representations of an ECAPA-TDNN model fine-tuned for speaker-verification, hence capturing speaker-related information such as timbre. The ECAPA-TDNN representations are mapped to a set of 32 hidden features using learned laten queries, which are then discretized using Finite Scalar Quantization Mentzer et al. (2023). Semantic tokens are derived from representations of the 11th, 14th and 16th layers of Wav2Vec 2.0 (XLSR-53) Pasad et al. (2023). These features are then quantized using single-codebook vector quantization, using factorized codes as done in DAC Kumar et al. (2023b). Tokens are converted back to raw audio using a Convolutional Neural Network (CNN) Decoder composed of ConvNeXt blocks. The language modeling framework is used to learn to generate semantic tokens, conditioned on the text-to-synthesis tokens and global tokens (which are prepended in the autoregressive sequence).
@@
is a non-autoregressive diffusion TTS system based on Flow Matching that does not rely on phonetic alignment nor duration predictor, operating on 24kHz sampling rate speech. During training, following the text-guided speech infilling task, speech mel spectrograms are first masked and then concatenated with text tokens. The resulting sequence is inputted to a Diffusion Transformer (DiT) Peebles and Xie (2023), which learns to generate the masked spectrogram region from gaussian noise using Flow Matching. During inference, the reference speech mel spectogram serves as the unmasked region, and is concatenated with the text to synthesize the input sequence. Sway sampling is used along with an Ordinary Differential Equation (ODE) solver to generate mel spectrograms that are converted back using Vocos Siuzdak (2024).
@@
is also a non-autoregressive TTS system, that follows a two-step masked generative modeling paradigm. It leverages two speech codec models to produce semantic and acoustinc discrete units from 24kHz sampling rate speech. Semantic tokens are derived from representations of the 17th layer of W2v-BERT 2.0, that are converted to discrete units using a single codebook of size 8,192 via classic Vector Quantization (VQ). Acoustic tokens are produced following RVQGAN Kumar et al. (2023a), and 8 codebooks of size 8,192 are used via Residual VQ (RVQ). To generate raw audio, MaskGCT follows a two-step process: first, it uses a Transformer to generate semantic tokens by iteratively predicting masked tokens conditioned on the text to synthesize, prompt tokens and previously generated semantic tokens ; second acoustic tokens are generated following a similar iterative procedure, but using an architecture similar to SoundStorm Borsos et al. (2023), and the generative process is conditioned with the prompt acoustic tokens and previously predicted semantic tokens. Finally, predicted acoustic tokens are converted back to raw audio using Vocos. Additionally, MaskGCT employs a Flow-Matching duration predictor to predict the target speech duration during inference.
IndexTTS2
is a recent autoregressive TTS system that enables fine-grained control over duration and emotions. It incorporates three modules: a Text-to-Semantic (T2S) module, a Semantic-to-Mel (S2M) module and a vocoder. The T2S module autoregressively generates semantic tokens derived from a dedicated Neural Speech Tokenizer, trained with the standard language modeling task, while the S2M module is trained to generate faithful mel-spectrogram from semantic tokens, trained with flow matching. The vocoder converts mel-spectrogram back into raw audio. The core innovation of IndexTTS2 relies on its capacity to control the duration and emotion of synthesized speech. In particular, a soft-instruction module derived from Qwen3 predicts an emotion distribution from the text description, that is used to condition the speech synthesis. This decoupled approach allows innovative disentanglement between emotional expression and speaker identity – an approach particularly suited for the task of synthesizing character quotation – leading to large improvements in expressive TTS.
Appendix G Subjective Experiments
@@
For the CMOS experiments, we ask the participants the following question: How expressive is this recording? Please focus on aspects of style (timbre, emotion and prosody), and ignore the aspects of content, grammar, or audio quality. For each instance, we give participants two audios containing one human sample and one generated sample, in a randomized order. We ask the participants to rate if the first audio sounds more expressive than the second on a scale.
For the MOS experiments, we ask the participants the following question: How natural (i.e human-sounding) is this recording? Please focus on examining the audio quality and naturalness, and ignore aspects of style (timbre, emotion, prosody). For each instance, we give participants a single audio, and ask them to rate the naturalness of the sample on an 1-5 Likert scale with 0.5 intervals.
Appendix H Investigating Contextual Matching
We showed that IndexTTS2 is able to produce high-quality expressive speech. However, we found during listening that some samples could exhibit a very strong emotion that is completely different from what one could expect from the context. To investigate the proportion of such cases, we leverage the reasoning traces produced by Gemini-2.5 Pro when providing a score for ContextMOS. These traces contain fine-grained details describing the expected emotion from the context, and what is conveyed appropriately or wrongly in the provided samples, sometimes with precise timestamps. To provide an overview of failure cases, we provide Gemini-2.5 Flash with a reasoning trace (in text format) only and prompt it to categorize each sample based on how its emotion matches the context. Figure 4 shows the result for each model and Figure 5 provides with the prompt used. While IndexTTS2 shows a large number of samples that convey the appropriate emotion, it also shows the largest number of samples with the wrong or opposite emotion. Providing contextual information to IndexTTS2 largely reduces the ratio of wrong/opposite emotions. Similarly, SparkTTS trained from scratch with contextual information that exhibits the largest ratio of samples with appropriate emotions across all variants of SparkTTS.
Overall, these results suggest that both models and human samples seem to insufficiently depict the appropriate emotions for a given quotation, showcasing challenges in faithfully impersonating characters during book reading.
Indeed, LibriVox recordings are recorded by amateur audiobook narrators, thus not all utterances are recorded with the same effort to deliver a perfect reading, which is the reason why ContextMOS ground-truth scores are not considerably higher. We believe that replicating our experiments with professional audiobook narrators samples would likely exhibit higher ContextMOS, drawing a clearer gap between human and synthesized quotations.
Appendix I Hardware Information
@@
You will be given a paragraph that includes reasoning traces of a previous audio analysis. This analysis was performed by a LLM that was tasked to judge wheter a speech sample of an audiobook utterance was appropriately delivered in terms of prosody and emotion, given the book context in which the utterance occurs.
Given the reasoning trace and the assessed score on a scale from 1 to 5, where 1 indicates complete failure at delivering the appropriate emotion, and 5 indicates complete success, your goal is to **classify** the prediction in **4 error categories**:
**Error Categories:**
1. Success: the speech sample successfully delivered the right prosody and emotion.
2. Opposite: the speech sample delivered natural prosody and emotion, but it was the opposite emotion/prosody that was expected given the context.
3. Insufficient: the speech sample delivered insufficient prosody and emotion given the context.
4. Wrong: the speech sample delivered natural prosody and emotion, but it was not the one expected given the context.
Now, you will be provided the reasoning trace, that include the prediction reasoning in natural language, and the integer score in a 1-5 scale.
**Reasoning Trace:**
{{{reasoning_trace}}}
**output_format**
You will output a json dictionary as follows:
"error_reasoning": str = Reasoning chain to explain why the prediction should be either one of the **Error Categories**.
"prediction": int = Your prediction between 1 and 4 based on the **Error Categories**.
- Note: Ensure the json structure is followed and the json output **MUST** be parsable without errors.(For example, escape the quotes wherever you add them inside a field of the json, all brackets and braces should be correctly paired.)
@@
"whispered": {
"id": "2",
"type": "verb",
"confidence": 10
"slowly": {
"id": "2",
"type": "adverb",
"confidence": 10
@@
"went": {
"id": "0",
"type": "verb",
"confidence": 9
"laughing": {
"id": "0",
"type": "verb",
"confidence": 9
@@
Your goal is to judge two audio samples that deliver the same text and analyze if a speech sample seems better than the other one on a particular **evaluation dimension** and determine the winner based on the scoring criterion.
You will rate each sample a score between 0 and 3 based on how well the speech corresponding to the target quotation within a particular text called **contextual_text** is delivered, then do their comparative analysis and provide your final judgement.
A sample should sound realistic and human-like, and should capture the specific nuances of the text.
You will be provided with the **contextual_text** which is the full text containing the quotation delivered in the speech sample and other side information.
the **text_category** and the **evaluation_criterion** corresponding to the **text_category**, in which you will be made aware of the **evaluation dimension** you will focus on, and the **scoring criteria** you will use to score the samples.
You will also be provided with the **output_format**, which dictates the format of the output you need to follow as a judger.
Finally, you will first be provided with the speech sample 1 **sample_1** and then speech sample 2 **sample_2**.
**contextual_text**
{{{context}}}
**text_category**
Emotions
**evaluation_criterion**
**Evaluation Dimension:**
- In this category, we want the expression of natural emotions and contextual voice acting, using variations in pitch, loudness, rhythm, etc.
- The sample contain speech delivered from a quotation present in the **contextual_text** and identified with <quote_start> and <quote_end> markers. Delivering a good quality speech means showing natural and strong expressiveness for the quoted dialogue, which aligns with the contextual information.
**Example:**
{{{Examples}}}
**Rating Scale:**
1: Fails to express emotions or contextual voice acting (whispering etc..).
2: The rendered emotions or contextual voice acting are not very natural.
3: Synthesises the quotation with natural emotions or contextual voice acting.
**Note:**
- The **contextual_text** will not explicitly state the emotion for the quotation, you have to infer that from the context.
- Samples demonstrating exaggerated expressiveness should not be rewarded more **UNLESS** the expressiveness features are relevant to the delivery of appropriated emotion and/or voice acting.
**Reasoning guidelines:**
{{{reasoning_guidelines}}}
NOTE: If the speech is very poor and does not synthesise the text correctly, you will provide a score of 0.
GLOBAL CONSIDERATIONS(**VERY IMPORTANT FOR COMPARISON**):
- It is imperative to compare the two samples **ONLY** on the basis of the **evaluation_dimension**, that means, you **WILL NOT** let the following types of **BIASES** affect your judgement:
{{{bias_mitigation}}}
**output_format**
You will output a json dictionary as follows:
"reasoning_sample_1": str = …
"reasoning_sample_2": str = …
"comparison": str = …
"score_1": int = …
"score_2": int = …
"winner": int = …
Now you will be provided with the speech from sample 1, please analyze it carefully.
**sample_1**
{{{audio_1}}}
Now you will be provided with the speech from sample 2, please analyze it carefully. After that provide the judgment following the **output_format** ensuring parsability.
**sample_2**
{{{audio_2}}}
Your goal is judge a single audio sample that deliver a specific quotation text. You will be provided with the **contextual_text** which is the full text containing the quotation delivered in the speech sample and other side information.
The target quotation that is spoken in the speech sample will be identified with <quote_start> and <quote_end> markers.
Given the **contextual_text** and the target quotation, you will rate the sample on a score between 0 and 5, based on how well the corresponding speech sample good quality speech means shows natural and strong expressiveness for the quoted dialogue, which aligns with the contextual information.
In particular you will focus on aspects of emotions and voice acting and ignore any other factors.
You will also be provided with the **output_format**, which dictates the format of the output you need to follow as a judger.
Finally, you will be provided with the speech sample **sample**.
**contextual_text**
{{{context}}}
**evaluation_criterion**
- We want to evaluate the expression of natural emotions and contextual voice acting, using variations in pitch, loudness, rhythm, etc.
**Example:**
{{{examples}}}
**Rating Scale:**
1: Fails completely to express emotions or contextual voice acting (whispering etc..).
2: Delivers an emotion or contextual voice acting that does not match the contextual information.
3: The rendered emotions or contextual voice acting are not very natural.
4: The rendered emotions or contextual voice acting is matching the contextual information but is not strong enough.
5: The delivered speech contains very natural emotions or contextual voice acting that perfectly align with the contextual information.
NOTE: If the speech is very poor and does not match the text correctly, you will provide a score of 0. GLOBAL CONSIDERATIONS(**VERY IMPORTANT FOR COMPARISON**):
- It is imperative to judge the sample **ONLY** on the basis of the **evaluation_criterion**, that means, you **WILL NOT** let the following types of **BIASES** affect your judgement:
- The acoustical quality of the audio, background noise or clarity.
- The gender and timbre features of the speaker. - Any other factors that are not related to the **evaluation_criterion**.
- Samples demonstrating exaggerated expressiveness should not be rewarded more **UNLESS** those features are relevant to the **evaluation_criterion**
**Reasoning guidelines:**
{{{reasoning_guidelines}}}
NOTE: If the speech is very poor and does not synthesise the text correctly, you will provide a score of 0.
GLOBAL CONSIDERATIONS(**VERY IMPORTANT FOR COMPARISON**):
- It is imperative to compare the two samples **ONLY** on the basis of the **evaluation_dimension**, that means, you **WILL NOT** let the following types of **BIASES** affect your judgement:
{{{bias_mitigation}}}
**output_format**
You will output a json dictionary as follows:
"reasoning": str = Reasoning chain based on the **Reasoning guidelines:**.
"score": int = Your score for the speech sample between 0 and 5, based on the **evaluation_criterion** and what you have mentioned in "reasoning".
Now you will be provided with the speech sample, please analyze it carefully.
**sample**
{{{audio}}}